montey:

Solve for x: 5x + 15 = 28

Aubree:

First do 5x = 28 - 15

Convert:

5x + 15 = 28 - 15 -15 5x=13 5x/5=13/5 x=13/5 is the furthest you can simplify

Aubree:

Then do x = answer / 5

supie:

\[5x + 15 = 28\] Now we must subtract 15 from both sides because its '+15' so we have to work backwards and subtract \[5x+15−15=28−15 \ so \ 5x=13 \] Now, we have to divide, so we have to add fraction bars As, you would know, a fraction bar is a division symbol so, _5x__ = _13____ 5 5 So \[x=blank\]